# Structural evolution of quiescent galaxies at intermediate redshift at UV and red rest-frame wavelengths

### v1

@@ 5.2.2 Sérsic index-stellar mass relation
The scaling relation between Sérsic index and stellar mass ($n-M_{\star}$ relation) probes the effects that the processes driving the mass assembly of quiescent systems have on the central concentration of their light profiles. Small galaxy samples and significant uncertainties in individual measurements of $n$ limit the number of studies that address the $n-M_{\star}$ relation (e.g., van Dokkum et al. 2010; Huertas-Company et al. 2013). These studies cover a broad redshift interval ($20<z<$ 2) and find heterogeneous trends with redshift. The results in the literature include both a decrease and a slight increase of $n$ with increasing redshift.
@@
With the dense HectoMAP sample we explore the evolution in $n-M_{\star}$ relation as a function of the rest-frame wavelength used for the estimates of structural parameters. We also explore the dependence of this relation on the the global stellar population age that separates newcomers from the aging subpopulation.
Figure 11: Same as the upper and middle panels of Figure 9, but for the Sérsic index.
@@
Figure 11 shows the $n-M_{\star}$ relation for HectoMAP quiescent galaxies in two different rest-frames. In each panel, black points and red crosses show the median sizes in 0.25 dex-wide bins of stellar mass in the UV and red rest-frame, respectively. We fit the relation $n^{RF}=n_{0}^{RF}+\gamma\log{(M_{\star}/10^{11}\text{M}_{\odot})}$ to the median values. The black solid and red dashed lines show the results in UV and red rest frame, respectively.
The slope of the $n-M_{\star}$ relation increases with redshift in both rest-frames. From $z\sim 0.25$ to $z\sim 0.55$, the slope increases by $\sim 63\%$ and $\sim 85\%$ in the UV and red rest-frames, respectively.
@@
The Sérsic indices in the UV are systematically smaller than those in red; the offset increases with increasing redshift. The zero point of the $n-M_{\star}$ relation $n_{0}^{RF}$, the characteristic Sérsic index of a $10^{11}$ quiescent galaxy, shows stronger evolution in the red than in the UV rest-frame. In the UV rest-frame, the zero point increases by $\sim 10\%$ (around $n_{0}^{\rm UV}\sim 3$) as the redshift increases from $z\sim 0.25$ to $z\sim 0.55$. In the red rest-frame the zero point increases from $n_{0}^{\rm red}=3.93\pm 0.05$ to $n_{0}^{\rm red}=4.81\pm 0.06$ ($\sim 22\%$) over the same redshift range.
Figure 12: Same as Figure 10, but for Sérsic index.
@@
We use interpolated values for the Sérsic index at $M_{\star}=10^{11}\,M_{\odot}$, based on the medians of measured $n$ values in the adjacent mass bins, as the characteristic Sérsic indices $n_{0}^{\mathrm{UV}}$ and $n_{0}^{\mathrm{red}}$ in two rest frames and show their evolution for the complete quiescent sample in Figure 12. The black circles with error bars display the redshift evolution of the characteristic Sérsic index and its scatter in the UV (upper panel) and red (lower panel) rest-frame. The black lines show the fit of the characteristic indices to the relation $n_{0}^{RF}(z)=a(1+z)^{b}$. The exponent of the relation in the red rest-frame is $b=+0.71\pm 0.13$, significantly steeper than the exponent in the UV ($b=+0.38\pm 0.08$).
@@
We also explore the redshift evolution of the characteristic Sérsic index at two rest-frame wavelengths as a function of the average $D_{n}4000$ for the populations of newcomers and aging galaxies (Sect. 4.2, color-coded points in Figure 12). In the UV rest-frame $M_{\star}\sim 10^{11}\,M_{\odot}$ newcomers are less centrally concentrated (i.e., more disk-like with $n_{0}^{\rm UV}\sim 2$) than the spheroidal ($n_{0}^{\rm UV}\sim 4$) aging quiescent galaxies. The two subpopulations show essentially constant $n_{0}^{\rm UV}(z)$. The evolution in the Sérsic index measured in rest-frame UV light for newcomers and aging galaxies is consistent with the redshift trend for the total sample.
@@
In the red rest-frame, the characteristic index $n_{0}^{\mathrm{red}}$ for both newcomers and aging galaxies increases with increasing redshift, albeit at different rates. For the aging subpopulation, the central light concentration increases negligibly with redshift. In contrast, the drop in the central concentration of newcomers is significant. At $z\sim 0.25$, newcomers have $\sim 2$ times smaller $n$ than aging galaxies; at $z\sim 0.55$, the two subpopulations exhibit similar $n$ values.
@@
A relatively constant value of $n$ for $M_{\star}\sim 10^{11}\,M_{\odot}$ quiescent galaxies in the rest-frame UV complements the lack of evolution in size $R_{0}^{\mathrm{UV}}$ of quiescent galaxy light profiles (Figure 10). The offset in the rest-frame UV values of $n$ between newcomers and the aging population illustrates the difference in the concentration of the blue light profiles. For newcomers, these light profiles are are disk-like. In contrast, the blue light regions of aging quiescent galaxies are spheroidal halos.
@@
For the subpopulation of newcomers characteristic Sérsic index increases substantially with redshift in the red rest-frame, from $n\sim 2.5$ at $z\sim 0.25$ to $n\sim 4.5$ at $z\sim 0.55$. This evolutionary trend implies that the distribution of the bulk of stellar mass in star-forming galaxies at $0.70.6<z<$ 0.7 (progenitors of newcomers at $z\sim 0.55$) is much more centrally concentrated than at $\Delta z\sim 0.2$ lower redshift.
@@
The scaling relation between Sérsic index $n$ and galaxy stellar mass in the UV and red rest-frames shows that all quiescent systems have blue light profiles that are less centrally concentrated than their light profiles at redder wavelengths. When separated into subpopulations of recently quenched systems (newcomers) and the oldest resident (aging) population based on their $D_{n}4000$ index, UV rest-frame profiles of young quiescent galaxies are, on average, disk-like, and the profiles of aging population in the same wavelength regime are spheroidal. In contrast with the size evolution in the red rest-frame (Section 5.2.1), the evolution in Sérsic index is very mild and, within the uncertainties, indicates little or no variation in the light profiles between $z\sim 0.55$ and $z\sim 0.25$ in both rest-frames. Together with larger $n_{0}^{RF}$ for the aging subpopulation in both rest-frames, the relatively constant central concentration of light profiles at $0.60.2<z<$ 0.6 points at the residual (and/or preceding) star formation (newcomers) and minor mergers (aging subpopulation) as the main drivers of the observed structural evolution of quiescent systems at intermediate redshifts.
@@ 6 Conclusion
We fit a single Sérsic model to the five-band images of the quiescent galaxy sample (Sect. 3.3), and use the best-fit linear relation between structural parameters and the wavelength of corresponding images to measure half-light radius and Sérsic index in two rest-frames: $3500$ (UV) and $7000$ (red). For each galaxy we measure the ratio between the logarithmic difference in the circularized half-light radius or Sérsic index and the logarithmic difference in the two rest-frame wavelengths of the two measurements (Sect. 4.1).
The dense HectoMAP spectro-photometric sample enables exploration of the evolution of the change in structural parameters with the rest-frame wavelength over four equally-spaced redshift bins covering the range $0.60.2<z<$ 0.6 (Sect. 4.1). Galaxy size decreases with rest-frame wavelength, independent of mass. The strength of this anti-correlation grows weakly with redshift. In contrast, the Sérsic index increases with rest-frame wavelength. The amplitude of this increase is essentially independent of mass and redshift. The distribution of the bulk of stellar mass in quiescent systems (traced by redder rest-frame wavelength) is more centrally concentrated and less extended than the distribution of younger and/or metal-poor stellar populations dominating galaxy light profiles in bluer rest-frame wavelengths. At intermediate redshift, the HectoMAP survey provides the measurements of structural parameters that cover a broad rest-frame wavelength range, previously available for statistically large spectroscopic samples only at $z<0.2$ (e.g., Kelvin et al. 2012; Kennedy et al. 2016).
@@
We study the size-stellar mass and Sérsic index-stellar mass relation for quiescent population in the UV and red rest-frames (Sects. 5.2.1 and 5.2.2). In the red rest-frame, the characteristic size of $10^{11}$ quiescent galaxies increases with decreasing redshift as $R_{0}^{\mathrm{red}}\propto(1+z)^{-0.98\pm 0.11}$, or $\sim 30\%$ from $z\sim 0.55$ to $z\sim 0.25$. This result broadly agrees with previous investigations based mostly on size measurement in a single (red) observed or rest-frame photometric band (van der Wel et al. 2014; Mowla et al. 2019; Damjanov et al. 2023). On the other hand, in the rest-frame UV, the change in the characteristic size is much slower and consistent with no evolution. As the redshift increases, the typical Sérsic index of a $10^{11}$ quiescent galaxy increases faster at longer wavelengths: it increases (modestly) by $\sim 7\%$ and $\sim 14\%$ in the UV and red rest-frames, respectively, from $z\sim 0.2$ to $z\sim 0.6$.
@@
The newcomers and aging subpopulation both show nearly constant (rest-frame) UV-based characteristic size and Sérsic index over the redshift range $0.60.2<z<$ 0.6, with the newcomers displaying an offset towards slightly ($\sim 20\%$) larger size and two times lower Sérsic index. In the red rest-frame, the characteristic sizes of both subpopulations increase at the same rate (within the uncertainties) as the size of a $M_{\star}=10^{11}\,M_{\odot}$ galaxy representing the complete quiescent sample. Conversely, a decrease in the Sérsic index with cosmic time is negligible for the resident population and significant for the newcomers ($\sim 42\%$ from $z\sim 0.6$ to $z\sim 0.2$). This difference arises because a fixed spectrograph fiber size, combined with strong color gradients and size evolution in newcomers, preferentially selects younger newcomers at lower redshifts (Section 5.2.2).
@@
Minor mergers are widely regarded as the main channel for growth of individual quiescent galaxies (e.g., Naab et al. 2009; Nipoti et al. 2009; Newman et al. 2012; Faisst et al. 2017; Hamadouche et al. 2022; Damjanov et al. 2024). In this context, the absence of change in the structural properties of the aging population in the rest-frame UV suggests that the star forming/low metallicity material is deposited into the extended regions of galaxy light profiles through minor mergers. As the merger frequency decreases with cosmic time, this material maintains a low level of residual star formation and settles into galaxy halos. On the other hand, size growth in the red rest-frame is driven by a combination of added stellar mass (through residual star formation) and a higher frequency of red satellites merging with massive quiescent galaxies at later epochs.
@@
The HectoMAP quiescent sample demonstrates the power of a large, complete mass-limited quiescent sample for exploring evolutionary trends in the extent and shape of light distributions, over a broad baseline of rest-frame wavelengths, for quiescent subpopulations segregated by average stellar population age. Ongoing and planned deep redshift surveys with unprecedentedly large area coverage (e.g., DESI Luminous Red Galaxies, PFS Galaxy Evolution Survey; Zhou et al. 2023; Greene et al. 2022) will enable the stellar population age-based selection of galaxy subpopulations in volume-limited samples reaching $z>1$. In parallel, space-based photometric surveys will provide robust multi-wavelength measurements of the half-light radii and Sérsic indices for large samples of quiescent galaxies over a a broad redshift range (e.g., Euclid; Euclid Collaboration et al. 2025b, a), enabling statistical probes of the changes in structural properties with the rest-frame wavelength at higher redshifts.
@@
The HectoMAP mass-complete spectro-photometric quiescent sample at intermediate redshift is a platform for the first statistical determination of galaxy structural parameters as a function of rest-frame wavelength at $0.60.2<z<$ 0.6. Spanning almost half of cosmic history, analysis of this sample yields critical insights into the dominant processes driving galaxy quenching and structural evolution during quiescence. We establish a foundational template for future surveys to extend these probes to even earlier epochs thus laying the groundwork for a more detailed, coherent picture of galaxy mass assembly.

### v2

@@ 5.2.2 Sérsic index-stellar mass relation
The scaling relation between Sérsic index and stellar mass ($n-M_{\star}$ relation) probes the effects that the processes driving the mass assembly of quiescent systems have on the central concentration of their light profiles. Small galaxy samples and significant uncertainties in individual measurements of $n$ limit the number of studies that address the $n-M_{\star}$ relation (e.g., van Dokkum et al. 2010; Huertas-Company et al. 2013). These studies cover a broad redshift interval ($0<z<2$) and find heterogeneous trends with redshift. The results in the literature include both a decrease and a slight increase of $n$ with increasing redshift. Most recent results based on large photometric samples from JWST show that the Sérsic index of quiescent galaxies with $\log(M_{\ast}/M_{\odot})>8$ slightly increases from $n\sim 4.5$ at $z\sim 2$ to $n\sim 5.4$ at $z\sim 0.5$ (Hamadouche et al. 2025).
@@
With the dense HectoMAP sample we explore the evolution in $n-M_{\star}$ relation as a function of the rest-frame wavelength used for the estimates of structural parameters. We also explore the dependence of this relation on the global stellar population age that separates newcomers from the aging subpopulation.
Figure 11: Sérsic index-stellar mass relation for quiescent galaxies in two rest frames. Analogously to Fig. 9, the black points and red crosses show the median galaxy concentration for the central values of equally spaced stellar mass bins in the two rest frames, UV and red, respectively, in four redshift bins with redshift increasing from left to right. The black solid and red dashed lines show the linear fit to the medians, and the thickness of dark gray and red bands around them corresponds to the uncertainties. The shaded bands represent the interquartile ranges of galaxy Sérsic indices across stellar mass bins.
@@
Figure 11 shows the $n-M_{\star}$ relation for HectoMAP quiescent galaxies in two different rest frames. In each panel, black points and red crosses show the median sizes in 0.25 dex-wide bins of stellar mass in the UV and red rest frame, respectively. We fit the relation $n^{RF}=n_{0}^{RF}+\gamma\log{(M_{\star}/10^{11}\text{M}_{\odot})}$ to the median values. The black solid and red dashed lines show the results in UV and red rest frame, respectively.
@@
The slope of the $n-M_{\star}$ relation increases with redshift in both rest frames. From $z\sim 0.25$ to $z\sim 0.55$, the slope increases by $\sim 63\%$ and $\sim 85\%$ in the UV and red rest frames, respectively. The Sérsic indices in the UV are systematically smaller than those in red; the offset increases with increasing redshift. The zero point of the $n-M_{\star}$ relation $n_{0}^{RF}$, the characteristic Sérsic index of a $10^{11}$M quiescent galaxy, shows stronger evolution in the red than in the UV rest-frame. In the UV rest-frame, the zero point increases by $\sim 10\%$ (around $n_{0}^{\rm UV}\sim 3$) as the redshift increases from $z\sim 0.25$ to $z\sim 0.55$. In the red rest frame the zero point increases from $n_{0}^{\rm red}=3.93\pm 0.05$ to $n_{0}^{\rm red}=4.81\pm 0.06$ ($\sim 22\%$) over the same redshift range.
Figure 12: Redshift evolution of the zero point of the Sérsic index-stellar mass relation (i.e., $n$ at the reference stellar mass of $10^{11}$ M) as a function of $D_{n}4000$. Analogously to Fig. 10, the top and bottom panels show the results in the UV and red rest frames, respectively. Black points show the median Sérsic index of the total sample of galaxies, and the sets of colder and warmer points show the zero points for the subpopulations of newcomers and aging galaxies, respectively. The error bars denote the interquartile range of the distribution. The dashed blue and dash-dotted orange lines are power law fits to the evolution of the zero point of the subpopulations of newcomers and aging galaxies, respectively.
@@
We use interpolated values for the Sérsic index at $M_{\star}=10^{11}\,M_{\odot}$, based on the medians of measured $n$ values in the adjacent mass bins, as the characteristic Sérsic indices $n_{0}^{\mathrm{UV}}$ and $n_{0}^{\mathrm{red}}$ in two rest frames and show their evolution for the complete quiescent sample in Fig. 12. The black circles with error bars display the redshift evolution of the characteristic Sérsic index and its scatter in the UV (upper panel) and red (lower panel) rest-frame. The black lines show the fit of the characteristic indices to the relation $n_{0}^{RF}(z)=a(1+z)^{b}$. The exponent of the relation in the red rest frame is $b=+0.71\pm 0.13$, significantly steeper than the exponent in the UV ($b=+0.38\pm 0.08$).
@@
We also explore the redshift evolution of the characteristic Sérsic index at two rest-frame wavelengths as a function of the average $D_{n}4000$ for the populations of newcomers and aging galaxies (Sect. 4.2, color-coded points in Fig. 12). In the UV rest-frame $M_{\star}\sim 10^{11}\,M_{\odot}$ newcomers are less centrally concentrated (i.e., more disk-like with $n_{0}^{\rm UV}\sim 2$) than the spheroidal ($n_{0}^{\rm UV}\sim 4$) aging quiescent galaxies. The two subpopulations show essentially constant $n_{0}^{\rm UV}(z)$. The evolution in the Sérsic index measured in rest-frame UV light for newcomers and aging galaxies is consistent with the redshift trend for the total sample.
@@
In the red rest frame, the characteristic index $n_{0}^{\mathrm{red}}$ for both newcomers and aging galaxies increases with increasing redshift, albeit at different rates. For the aging subpopulation, the central light concentration increases negligibly with redshift. In contrast, the drop in the central concentration of newcomers is significant. At $z\sim 0.25$, newcomers have $\sim 2$ times smaller $n$ than aging galaxies; at $z\sim 0.55$, the two subpopulations exhibit similar $n$ values. Equivalently to the characteristic size $R_{0}^{RF}$ (Sect. 5.2.1), the evolutionary trends in $n_{0}^{RF}$ at two rest-frame wavelengths for newcomers and aging population are not sensitive to the uncertainties in $D_{n}4000$ measurements.
@@
A relatively constant value of $n$ for $M_{\star}\sim 10^{11}\,M_{\odot}$ quiescent galaxies in the rest-frame UV complements the lack of evolution in size $R_{0}^{\mathrm{UV}}$ of quiescent galaxy light profiles (Fig. 10). The offset in the rest-frame UV values of $n$ between newcomers and the aging population illustrates the difference in the concentration of the blue light profiles. For newcomers, these light profiles are disk-like. In contrast, the blue light regions of aging quiescent galaxies are spheroidal halos.
@@
For the subpopulation of newcomers characteristic Sérsic index increases substantially with redshift in the red rest frame, from $n\sim 2.5$ at $z\sim 0.25$ to $n\sim 4.5$ at $z\sim 0.55$. This evolutionary trend implies that the distribution of the bulk of stellar mass in star-forming galaxies at $0.6<z<0.7$ (progenitors of newcomers at $z\sim 0.55$) is much more centrally concentrated than at $\Delta z\sim 0.2$ lower redshift.
@@
The scaling relation between Sérsic index $n$ and galaxy stellar mass in the UV and red rest frames shows that all quiescent systems have blue light profiles that are less centrally concentrated than their light profiles at redder wavelengths. When separated into subpopulations of recently quenched systems (newcomers) and the oldest resident (aging) population based on their $D_{n}4000$ index, UV rest-frame profiles of young quiescent galaxies are, on average, disk-like, and the profiles of aging population in the same wavelength regime are spheroidal. In contrast with the size evolution in the red rest frame (Sect. 5.2.1), the evolution in Sérsic index is very mild and, within the uncertainties, indicates little or no variation in the light profiles between $z\sim 0.55$ and $z\sim 0.25$ in both rest frames. Together with larger $n_{0}^{RF}$ for the aging subpopulation in both rest frames, the relatively constant central concentration of light profiles at $0.2<z<0.6$ points at the residual (and/or preceding) star formation (newcomers) and minor mergers (aging subpopulation) as the main drivers of the observed structural evolution of quiescent systems at intermediate redshifts.
@@ 6 Conclusion
We fit a single Sérsic model to the five-band images of the quiescent galaxy sample (Sect. 3.3) and used the best-fit linear relation between structural parameters and the wavelength of corresponding images to measure half-light radius and Sérsic index in two rest frames: $3500$ Å (UV) and $7000$ Å (red). For each galaxy, we measured the ratio between the logarithmic difference in the circularized half-light radius or Sérsic index and the logarithmic difference in the two rest-frame wavelengths of the two measurements (Sect. 4.1).
The dense HectoMAP spectro-photometric sample enables an exploration of the evolution of the change in structural parameters with the rest-frame wavelength over four equally spaced redshift bins covering the range $0.2<z<0.6$ (Sect. 4.1). Galaxy size decreases with rest-frame wavelength, independent of mass. The strength of this anti-correlation grows weakly with redshift. In contrast, the Sérsic index increases with rest-frame wavelength. The amplitude of this increase is essentially independent of mass and redshift. The distribution of the bulk of stellar mass in quiescent systems (traced by redder rest-frame wavelength) is more centrally concentrated and less extended than the distribution of younger and/or metal-poor stellar populations dominating galaxy light profiles in bluer rest-frame wavelengths. At intermediate redshift, the HectoMAP survey provides the measurements of structural parameters that cover a broad rest-frame wavelength range, previously available for statistically large spectroscopic samples only at $z<0.2$ (e.g., Kelvin et al. 2012; Kennedy et al. 2016).
@@
We study the size-stellar mass and Sérsic index-stellar mass relation for quiescent population in the UV and red rest frames (Sects. 5.2.1 and 5.2.2). In the red rest frame, the characteristic size of $10^{11}$M quiescent galaxies increases with decreasing redshift as $R_{0}^{\mathrm{red}}\propto(1+z)^{-0.98\pm 0.11}$, or $\sim 30\%$ from $z\sim 0.55$ to $z\sim 0.25$. This result broadly agrees with previous investigations based mostly on size measurement in a single (red) observed or rest-frame photometric band (van der Wel et al. 2014; Mowla et al. 2019; Damjanov et al. 2023). On the other hand, in the rest-frame UV, the change in the characteristic size is much slower and consistent with no evolution. As the redshift increases, the typical Sérsic index of a $10^{11}$M quiescent galaxy increases faster at longer wavelengths: it increases (modestly) by $\sim 7\%$ and $\sim 14\%$ in the UV and red rest frames, respectively, from $z\sim 0.2$ to $z\sim 0.6$.
@@
The newcomers and aging subpopulation both show nearly constant (rest-frame) UV-based characteristic size and Sérsic index over the redshift range $0.2<z<0.6$, with the newcomers displaying an offset towards slightly ($\sim 20\%$) larger size and two times lower Sérsic index. In the red rest frame, the characteristic sizes of both subpopulations increase at the same rate (within the uncertainties) as the size of a $M_{\star}=10^{11}\,M_{\odot}$ galaxy representing the complete quiescent sample. Conversely, a decrease in the Sérsic index with cosmic time is negligible for the resident population and significant for the newcomers ($\sim 42\%$ from $z\sim 0.6$ to $z\sim 0.2$). This difference arises because a fixed spectrograph fiber size, combined with strong color gradients and size evolution in newcomers, preferentially selects younger newcomers at lower redshifts (Sect. 5.2.2).
@@
Minor mergers are widely regarded as the main channel for growth of individual quiescent galaxies (e.g., Naab et al. 2009; Nipoti et al. 2009; Newman et al. 2012; Faisst et al. 2017; Hamadouche et al. 2022; Damjanov et al. 2025; Hamadouche et al. 2025). In this context, the absence of change in the structural properties of the aging population in the rest-frame UV suggests that the star-forming-and-low-metallicity material is deposited into the extended regions of galaxy light profiles through minor mergers. As the merger frequency decreases with cosmic time, this material maintains a low level of residual star formation and settles into galaxy halos. On the other hand, size growth in the red rest frame is driven by a combination of added stellar mass (through residual star formation) and a higher frequency of red satellites merging with massive quiescent galaxies at later epochs.
@@
The HectoMAP quiescent sample demonstrates the power of a large, complete mass-limited quiescent sample for exploring evolutionary trends in the extent and shape of light distributions, over a broad baseline of rest-frame wavelengths, for quiescent subpopulations segregated by average stellar population age. Ongoing and planned deep redshift surveys with unprecedentedly large area coverage (e.g., DESI Luminous Red Galaxies, PFS Galaxy Evolution Survey; Zhou et al. 2023; Greene et al. 2022) will allow for stellar population age-based selection of galaxy subpopulations in volume-limited samples up to $z>1$. In parallel, space-based photometric surveys will provide robust multiwavelength measurements of the half-light radii and Sérsic indices for large samples of quiescent galaxies over a a broad redshift range (e.g., Euclid; Euclid Collaboration et al. 2025c, b), enabling statistical probes of the changes in structural properties with the rest-frame wavelength at higher redshifts.
@@
The HectoMAP mass-complete spectro-photometric quiescent sample at intermediate redshift is a platform for the first statistical determination of galaxy structural parameters as a function of the rest-frame wavelength at $0.2<z<0.6$. Spanning almost half of cosmic history, our analysis of this sample yields critical insights into the dominant processes driving galaxy quenching and structural evolution during quiescence. We have established a foundational template for future surveys to extend these probes to even earlier epochs, thereby laying the groundwork for a more detailed, coherent picture of galaxy mass assembly.
Data availability
Table 2 is only available in electronic form at the CDS via anonymous ftp to cdsarc.u-strasbg.fr (130.79.128.5) or via http://cdsweb.u-strasbg.fr/cgi-bin/qcat?J/A+A/.